OpenStudy (anonymous):

breakfast at the local cafe costs harry $38.70 without taxes. He wants to leave a 15 percent tip. how much should he tip?

OpenStudy (anonymous):

I'm not sure if you have to consider taxes, lets not. if you want to know x% of some number you multiply it with x/100 15% of 38.70 = 0.15 * 38.70 =

OpenStudy (anonymous):

im sorry but im still not understanding

OpenStudy (anonymous):

Well 15 percent means 15 hundredths. 15 hundredths * $38.70 = $5.8, the tip.